According to the following reaction, how many grams of aluminum sulfate will be formed upon the complete reaction of 22.5 grams of sulfuric acid with excess aluminum oxide?

aluminum oxide (s) + sulfuric acid (aq) aluminum sulfate (aq) + water (l)
1
Expert's answer
2020-03-24T13:56:41-0400

Al2O3 + 3H2SO4 = Al2(SO4)2 + 3H2O

n(H2SO4) = m/M = 22.5g/98g/mol = 0.23 mol

According to the equation:

n(Al2(SO4)2) = 0.23/3 = 0.0767 mol

m(Al2(SO4)2) = n*M = 0.0767mol*342g/mol = 26.23 g